Consumer Debt
11. Suppose you want to buy a car. The dealer offers a financial package consisting of 6% APR compounded monthly for a term of five years. Suppose that you want your monthly payments to be at most $320. What is the maximum amount you should finance? Give your answer to the nearest dollar.

Answer by addingup(3677)

5*12 = 60 5 years equals 60 months. And you make a monthly payment of $320:
320*60 = 19,200
:
PV = 19,200/[(1+(0.06/12))^(12*5)] = 14,234.35 is what you should finance